# Well-posedness and stability of the Lagrange representation of the n-D wave equation via boundary triples

### v1

@@ Abstract.
We study the Lagrange representation of the wave equation with generalized Laplacian $\operatorname{div}T\nabla$. We allow the coefficients—the Young modulus $T$ and the density $\rho$—to be $\mathrm{L}^{\infty}$ or even nonlocal operators. Moreover, the Lipschitz boundary of the domain $\Omega$ can be split into several parts admitting Dirichlet, Neumann and/or Robin-boundary conditions of displacement, velocity and stress. We show well-posedness of this classical model of the wave equation utilizing boundary triple theory for skew-adjoint operators. In addition we show semi-uniform stability of solutions under slightly stronger assumptions by means of a spectral result.
@@ 1. Introduction
$\frac{\partial}{\partial t}\begin{pmatrix}\nabla w\\ \rho w_{t}\end{pmatrix}=\underbrace{\begin{pmatrix}0&\nabla\\ \operatorname{div}&0\end{pmatrix}}_{\eqqcolon\mathrlap{\mathcal{J}_{1}}}\underbrace{\begin{pmatrix}T&0\\ 0&\frac{1}{\rho}\end{pmatrix}}_{\eqqcolon\mathrlap{\mathcal{Q}_{1}}}\begin{pmatrix}\nabla w\\ \rho w_{t}\end{pmatrix}.$
This is the so-called Dirac representation, which is comprised of the formally skew-adjoint $\mathcal{J}_{1}$ (representing the underlying Dirac structure/subspace) and the bounded, positive and self-adjoint $\mathcal{Q}_{1}$ (representing the Lagrangian structure/subspace). However in [5] an alternative so-called Lagrangian representation was proposed, which uses $\begin{pmatrix}w\\ \rho w_{t}\end{pmatrix}$ as state variable. The corresponding system is
$\frac{\partial}{\partial t}\begin{pmatrix}w\\ \rho w_{t}\end{pmatrix}=\underbrace{\begin{pmatrix}0&\mathrm{I}\\ -\mathrm{I}&0\end{pmatrix}}_{\eqqcolon\mathrlap{\mathcal{J}_{2}}}\underbrace{\begin{pmatrix}-\operatorname{div}T\nabla&0\\ 0&\frac{1}{\rho}\end{pmatrix}}_{\eqqcolon\mathrlap{\mathcal{Q}_{2}}}\begin{pmatrix}w\\ \rho w_{t}\end{pmatrix}.$ (2)
Similar to the Dirac representation, $\mathcal{J}_{2}$ is skew-adjoint and $\mathcal{Q}_{2}$ is formally self-adjoint and positive. In contrast to the Dirac-representation however, $\mathcal{J}_{2}$ is bounded and $\mathcal{Q}_{2}$ is unbounded. This formulation also corresponds to the standard method to turn the wave equation into a first order problem, cf. [17, Sec. 7.4].
Naturally, employing terms like (un)bounded and self-adjoint warrants specifying the state spaces. For the Dirac representation one chooses $\mathrm{L}^{2}(\Omega)^{d}\times\mathrm{L}^{2}(\Omega)$ and for the Lagrange representation it comes naturally to choose $\mathrm{H}^{1}(\Omega)\times\mathrm{L}^{2}(\Omega)$. In order to analyze well-posedness for the Dirac representation it suffices to analyze $\mathcal{J}_{1}$ as $\mathcal{Q}_{1}$ can be incorporated into the inner product and thus does not play a role, cf. [14, Lem. 7.2.3]. We want to mimic this approach for the Lagrange representation, but since $\mathcal{Q}_{2}$ is unbounded, it does not immediately provide an equivalent inner product on the state space, i.e., we want to regard the following almost inner product [^1] We are a bit sloppy here in not specifying the regularity of the state variables.
$\langle x,y\rangle_{\mathcal{Q}_{2}}\coloneqq\langle\mathcal{Q}_{2}x,y\rangle=\left\langle\begin{pmatrix}-\operatorname{div}T\nabla&0\\ 0&\frac{1}{\rho}\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix},\begin{pmatrix}y_{1}\\ y_{2}\end{pmatrix}\right\rangle\\ =\langle\tfrac{1}{\rho}x_{2},y_{2}\rangle-\langle\operatorname{div}T\nabla x_{1},y_{1}\rangle\text{.}$
@@
where $\nu$ denotes the unit normal vector on $\partial\Omega$. It is worth pointing out that this expression is only an inner product on the state space $\mathrm{H}^{1}(\Omega)\times\mathrm{L}^{2}(\Omega)$ if we additionally impose a boundary condition such as $\nu\cdot T\nabla x_{1}\big{|}_{\partial\Omega}=-k_{1}x_{1}\big{|}_{\partial\Omega}$, where $k_{1}$ is a positive semi-definite operator [^2] We are in particular interested in multiplication operators that may vanish on parts of the boundary. on $\mathrm{L}^{2}(\partial\Omega)$. In that case we obtain
@@
and we can appeal to the Friedrichs/Poincaré inequality. Alternatively, the boundary condition $x_{1}\big{|}_{\partial\Omega}=0=y_{1}\big{|}_{\partial\Omega}$ would eliminate the boundary parts completely. In either case, the above expression induces an equivalent inner product on the state space $\mathrm{H}^{1}(\Omega)\times\mathrm{L}^{2}(\Omega)$. Note that the energy of a state $x\in\mathrm{H}^{1}(\Omega)\times\mathrm{L}^{2}(\Omega)$ is given by
$\mathrm{E}(x)\coloneqq\langle\tfrac{1}{\rho}x_{2},x_{2}\rangle+\langle T\nabla x_{1},\nabla x_{1}\rangle.$
Therefore the inner product $\langle\boldsymbol{\cdot},\boldsymbol{\cdot}\rangle_{\mathcal{Q}_{2}}$ is composed of an energy part and a boundary part. We will construct a boundary triple for the operator in the abstract Cauchy equation (2), which will enable us to parameterize all dissipative boundary conditions, i.e., boundary conditions where the solution does not grow. Loosely speaking we will come to the conclusion that boundary conditions of the form
$k_{1}x_{1}\big{|}_{\partial\Omega}+\nu\cdot T\nabla x_{1}\big{|}_{\partial\Omega}+k_{2}\tfrac{1}{\rho}x_{2}\big{|}_{\partial\Omega}=0$
@@
where $k_{1}$ is the positive semi-definite operator from before and $k_{2}$ is another positive semi-definite operator on $\mathrm{L}^{2}(\partial\Omega)$, will be well-posed. We point out that in contrast to the Dirac representation, cf. [15], we can formulate boundary conditions that involve the displacement $w$ itself (additionally to the velocity $\partial_{t}w$ and the normal stress $\nu\cdot T\nabla w$).
Additionally, since $k_{1}$ and $k_{2}$ are allowed to be semi-definite we can split the boundary into five (possibly empty) parts and allow the following boundary conditions:
$w&=0&\quad&\text{on}\ \Gamma_{0}\\ \nu\cdot T\nabla w&=0&&\text{on}\ \Gamma_{1}\\ k_{1}w+\nu\cdot T\nabla w&=0&&\text{on}\ \Gamma_{2}\\ \nu\cdot T\nabla w+k_{2}\partial_{t}w&=0&&\text{on}\ \Gamma_{3}\\ k_{1}w+\nu\cdot T\nabla w+k_{2}\partial_{t}w&=0&&\text{on}\ \Gamma_{4}\end{aligned}$ (3)
where we assume $k_{1}$ to be nonzero almost everywhere on $\Gamma_{2}\cup\Gamma_{4}$ and $k_{2}$ to be nonzero almost everywhere on $\Gamma_{3}\cup\Gamma_{4}$, i.e., the vanishing parts are being accounted for by $\Gamma_{0}$ and $\Gamma_{3}$ or by $\Gamma_{0},\Gamma_{1}$ and $\Gamma_{2}$ respectively.
Note that $\Gamma_{0}=\emptyset$ and $\Gamma_{2}\cup\Gamma_{4}=\emptyset$ ($k_{1}=0$) are allowed, but not simultaneously, i.e., we require
$k_{1}\neq 0\quad\text{or}\quad\Gamma_{0}\neq\emptyset.$
@@
We will pursue an alternative strategy, following [13] in our approach, settling for a weaker notion than exponential stability, by the name of “semi-uniform stability”, cf. Section 4. Slightly stronger assumptions on the coefficients $T$ and $\rho$ in conjunction with dissipativity will prove enough to show at least semi-uniform stability of solutions to Equation 1. In particular, we have to show that there is no spectrum on the imaginary axis. A similar approach has been applied to Maxwell’s equations in [21].
@@
    Let $T\in\mathcal{L}_{\mathrm{b}}(\mathrm{L}^{2}(\Omega;\mathbb{C}^{d}))$ with $c^{-1}\mathrm{I}<T<c\mathrm{I}$ for some $c>0$.
@@
    Let $w_{0}\in\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)$ and $w_{1}\in\mathrm{L}^{2}(\Omega)$. [^4] Here $\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)=\{v\in\mathrm{H}^{1}(\Omega)\,|\,\mathopen{}v|_{\Gamma_{0}}=0\}$. We will come back to this space in Section 3.
  5. (A5)
    The boundary is split into five open disjoint and possible empty parts $\Gamma_{0},\Gamma_{1},\Gamma_{2},\Gamma_{3},\Gamma_{4}\subseteq\partial\Omega$ that satisfy
@@
    For stability additionally: $\exists\Gamma\subseteq\partial\Omega\setminus\overline{\Gamma_{0}}$ non-empty open such that $\operatorname{supp}k_{2}\supseteq\Gamma$ [^5] This just means that the damping acts on an open set.
@@
    Let $a,b\in\mathrm{L}^{\infty}(\Omega)$. [^6] More general assumptions are possible, cf. sections 3 and 4.
@@
Assumption (A6) is not as general as possible, because otherwise the formulation would be a bit cumbersome. We could relax the conditions to: $k_{1},k_{2}\in\mathcal{L}_{\mathrm{b}}(\mathrm{L}^{2}(\partial\Omega\setminus\Gamma_{0}))$ positive semi-definite such that $1\notin\ker k_{1}$ or $\Gamma_{0}\neq\emptyset$. For stability we additionally ask for
$k_{2}f=0\quad\implies\quad\exists\Gamma\subseteq\partial\Omega\setminus\overline{\Gamma_{0}}\ \text{open}\ :f\big{|}_{\Gamma}=0.$ (4)
The condition $1\notin\ker k_{1}$ is necessary for Friedrichs/Poincaré’s inequality and (4) is necessary to apply the unique continuation principle.
In the following we first present a short preliminary section, which serves as a reminder for well-established concepts such as trace operators and boundary triples, which we will employ to prove our main results about well-posedness (Section 3) and stability (Section 4). We will end our article with a conclusion and a short appendix on the applicability of a unique continuation theorem in Section 4 and an additional result regarding regularity in the case of the trace maps taking values in $\mathrm{L}^{2}(\partial\Omega)$.
@@ 2. Preliminaries
This section is largely a short recapitulation of established theory that we present both for the convenience of the reader as well as to establish notation, that we are going to use in consecutive sections.
@@ 2.1. Sobolev Spaces
We clarify/introduce some notation first: We let $\mathrm{H}^{1}(\Omega)$ be the space of all $\mathrm{L}^{2}$-functions with distributional derivative also in $\mathrm{L}^{2}(\Omega)$ together with the norm $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}\coloneqq\sqrt{\lVert\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}^{2}+\lVert\nabla\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}^{2}}$. In other words, $\mathrm{H}^{1}(\Omega)=\bigl{(}\operatorname{dom}(\nabla),\lVert\boldsymbol{\cdot}\rVert_{\operatorname{dom}(\nabla)}\bigr{)}$, where $\nabla$ is the weak gradient on $\mathrm{L}^{2}(\Omega)$. Similarly, the corresponding space for the divergence operator $\operatorname{div}f\coloneqq\sum_{i=1}^{d}\partial_{i}f_{i}$ is

### v2

@@ Abstract.
We study the Lagrange representation of the wave equation with generalized Laplacian $\operatorname{div}T\nabla$. We allow the coefficients—the Young modulus $T$ and the density $\rho$—to be $\mathrm{L}^{\infty}$ or even nonlocal operators. Moreover, the Lipschitz boundary of the domain $\Omega$ can be split into several parts admitting Dirichlet, Neumann and/or Robin-boundary conditions of displacement, velocity and stress. We show well-posedness of this classical model of the wave equation utilizing boundary triple theory for skew-adjoint operators. In addition, we show semi-uniform stability of solutions under slightly stronger assumptions by means of a spectral result.
@@
Supported by the state of Saxony via a graduate student stipend
@@ 1. Introduction
$\frac{\partial}{\partial t}\begin{pmatrix}\nabla w\\ \rho w_{t}\end{pmatrix}=\underbrace{\begin{pmatrix}0&\nabla\\ \operatorname{div}&0\end{pmatrix}}_{\eqcolon\mathrlap{\mathcal{J}_{1}}}\underbrace{\begin{pmatrix}T&0\\ 0&\frac{1}{\rho}\end{pmatrix}}_{\eqcolon\mathrlap{\mathcal{Q}_{1}}}\begin{pmatrix}\nabla w\\ \rho w_{t}\end{pmatrix}$.
This is the so-called Dirac representation, which is comprised of the formally skew-adjoint $\mathcal{J}_{1}$ (representing the underlying Dirac structure/subspace) and the bounded, positive and self-adjoint $\mathcal{Q}_{1}$ (representing the Lagrangian structure/subspace). However, in [7] an alternative so-called Lagrange representation was proposed, which uses $\begin{pmatrix}w\\ \rho w_{t}\end{pmatrix}$ as state variable. The corresponding system is
$\frac{\partial}{\partial t}\begin{pmatrix}w\\ \rho w_{t}\end{pmatrix}=\underbrace{\begin{pmatrix}0&\mathrm{I}\\ -\mathrm{I}&0\end{pmatrix}}_{\eqcolon\mathrlap{\mathcal{J}_{2}}}\underbrace{\begin{pmatrix}-\operatorname{div}T\nabla&0\\ 0&\frac{1}{\rho}\end{pmatrix}}_{\eqcolon\mathrlap{\mathcal{Q}_{2}}}\begin{pmatrix}w\\ \rho w_{t}\end{pmatrix}$. (2)
Similar to the Dirac representation, $\mathcal{J}_{2}$ is skew-adjoint and $\mathcal{Q}_{2}$ is formally self-adjoint and positive. In contrast to the Dirac-representation however, $\mathcal{J}_{2}$ is bounded and $\mathcal{Q}_{2}$ is unbounded. This formulation also corresponds to the standard method to turn the wave equation into a first order problem, cf. [21, Sec. 7.4].
Naturally, employing terms like (un)bounded and self-adjoint warrants specifying the state spaces. For the Dirac representation one chooses $\mathrm{L}^{2}(\Omega;\mathbb{C}^{d})\times\mathrm{L}^{2}(\Omega;\mathbb{C})$ and for the Lagrange representation it comes naturally to choose $\mathrm{H}^{1}(\Omega;\mathbb{C})\times\mathrm{L}^{2}(\Omega;\mathbb{C})$. In order to analyze well-posedness for the Dirac representation, it suffices to analyze $\mathcal{J}_{1}$, as $\mathcal{Q}_{1}$ can be incorporated into the inner product and thus does not play a role, cf. [17, Lem. 7.2.3]. We want to mimic this approach for the Lagrange representation, but since $\mathcal{Q}_{2}$ is unbounded, it does not immediately provide an equivalent inner product on the state space, i.e., we want to regard the following almost inner product
$\langle x,y\rangle_{\mathcal{Q}_{2}}\coloneq\langle\mathcal{Q}_{2}x,y\rangle=\left\langle\begin{pmatrix}-\operatorname{div}T\nabla&0\\ 0&\frac{1}{\rho}\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix},\begin{pmatrix}y_{1}\\ y_{2}\end{pmatrix}\right\rangle\\ =\langle\tfrac{1}{\rho}x_{2},y_{2}\rangle-\langle\operatorname{div}T\nabla x_{1},y_{1}\rangle\text{.}$
@@
where $\nu$ denotes the unit normal vector on $\partial\Omega$. It is worth pointing out that this expression is only an inner product on the state space $\mathrm{H}^{1}(\Omega;\mathbb{C})\times\mathrm{L}^{2}(\Omega;\mathbb{C})$ if we additionally impose a boundary condition such as $\nu\cdot T\nabla x_{1}\big|_{\partial\Omega}=-k_{1}x_{1}\big|_{\partial\Omega}$, where $k_{1}$ is a positive semi-definite operator [^1] We are in particular interested in multiplication operators that may vanish on parts of the boundary. on $\mathrm{L}^{2}(\partial\Omega;\mathbb{C})$. In that case we obtain
@@
and we can appeal to the Friedrichs/Poincaré inequality. Alternatively, the boundary condition $x_{1}\big|_{\partial\Omega}=0=y_{1}\big|_{\partial\Omega}$ would eliminate the boundary parts completely. In either case, the above expression induces an equivalent inner product on the state space $\mathrm{H}^{1}(\Omega;\mathbb{C})\times\mathrm{L}^{2}(\Omega;\mathbb{C})$. Note that the energy of a state $x=\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\in\mathrm{H}^{1}(\Omega;\mathbb{C})\times\mathrm{L}^{2}(\Omega;\mathbb{C})$ is given by
$\mathrm{E}(x)\coloneq\langle\tfrac{1}{\rho}x_{2},x_{2}\rangle+\langle T\nabla x_{1},\nabla x_{1}\rangle.$
Therefore the inner product $\langle\boldsymbol{\cdot},\boldsymbol{\cdot}\rangle_{\mathcal{Q}_{2}}$ is composed of an energy part and a boundary part. We will construct a boundary triple for the operator in the abstract Cauchy equation (2), which will enable us to parameterize all dissipative boundary conditions, i.e., boundary conditions where the solution does not grow. Loosely speaking, we will come to the conclusion that boundary conditions of the form
$k_{1}x_{1}\big|_{\partial\Omega}+\nu\cdot T\nabla x_{1}\big|_{\partial\Omega}+k_{2}\tfrac{1}{\rho}x_{2}\big|_{\partial\Omega}=0,$
@@
where $k_{1}$ is the positive semi-definite operator from before and $k_{2}$ is another positive semi-definite operator on $\mathrm{L}^{2}(\partial\Omega;\mathbb{C})$, will be well-posed. We point out that in contrast to the Dirac representation, cf. [18, 2], we can formulate boundary conditions that involve the displacement $w$ itself (in addition to the velocity $\partial_{t}w$ and the normal stress $\nu\cdot T\nabla w$).
Additionally, since $k_{1}$ and $k_{2}$ are allowed to be semi-definite, we can split the boundary into five (possibly empty) parts and allow the following boundary conditions:
$\begin{aligned} w&=0&\quad&\text{on}\ \Gamma_{0},\\ \nu\cdot T\nabla w&=0&&\text{on}\ \Gamma_{1},\\ k_{1}w+\nu\cdot T\nabla w&=0&&\text{on}\ \Gamma_{2},\\ \nu\cdot T\nabla w+k_{2}\partial_{t}w&=0&&\text{on}\ \Gamma_{3},\\ k_{1}w+\nu\cdot T\nabla w+k_{2}\partial_{t}w&=0&&\text{on}\ \Gamma_{4},\end{aligned}$ (3)
where we assume $k_{1}$ to be nonzero almost everywhere on $\Gamma_{2}\cup\Gamma_{4}$ and $k_{2}$ to be nonzero almost everywhere on $\Gamma_{3}\cup\Gamma_{4}$, i.e., the vanishing parts are being accounted for by $\Gamma_{0},\Gamma_{1}$ and $\Gamma_{3}$ or by $\Gamma_{0},\Gamma_{1}$ and $\Gamma_{2}$ respectively.
Note that $\Gamma_{0}=\emptyset$ or $\Gamma_{2}\cup\Gamma_{4}=\emptyset$ ($k_{1}=0$) is allowed, but not simultaneously, i.e., we require
$\Gamma_{0}\neq\emptyset\quad\text{or}\quad k_{1}\neq 0.$
@@
We will pursue an alternative strategy, following [16] in our approach, settling for a weaker notion than exponential stability, by the name of “semi-uniform stability”, cf. Section 4. Slightly stronger assumptions on the coefficients $T$ and $\rho$ in conjunction with dissipativity will prove enough to show at least semi-uniform stability of solutions to problem (1). In particular, we have to show that there is no spectrum on the imaginary axis. A similar approach has been applied to Maxwell’s equations in [26].
@@
[…]
@@
    The boundary is split into five open, disjoint and possible empty parts $\Gamma_{0},\Gamma_{1},\Gamma_{2},\Gamma_{3},\Gamma_{4}\subseteq\partial\Omega$ that satisfy
@@
    Let $w_{0}\in\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega;\mathbb{C})$ and $w_{1}\in\mathrm{L}^{2}(\Omega;\mathbb{C})$. [^4] Here $\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)=\{v\in\mathrm{H}^{1}(\Omega)\,|\,\mathopen{}v|_{\Gamma_{0}}=0\}$. We will come back to this space in Section 3.
  6. (A6)
@@
    For stability additionally: $\exists\Gamma\subseteq\partial\Omega\setminus\overline{\Gamma_{0}}$ non-empty open such that $\operatorname{supp}k_{2}\supseteq\Gamma$. [^5] This just means that the damping acts on an open set.
@@
    Let $a,b\in\mathrm{L}^{\infty}(\Omega;\mathbb{C})$. [^6] More general assumptions are possible, cf. sections 3 and 4.
@@
Assumption (A6) is not as general as possible, because otherwise the formulation would be a bit cumbersome. We could relax the conditions to: $k_{1},k_{2}\in\mathcal{L}_{\mathrm{b}}(\mathrm{L}^{2}(\partial\Omega\setminus\overline{\Gamma_{0}}))$ positive semi-definite such that $1\notin\ker k_{1}$ or $\Gamma_{0}\neq\emptyset$. For stability we additionally ask for
$k_{2}f=0\quad\implies\quad\exists\Gamma\subseteq\partial\Omega\setminus\overline{\Gamma_{0}}\ \text{open and non-empty}\,:\,f\big|_{\Gamma}=0.$ (4)
The condition $1\notin\ker k_{1}$ is necessary for Friedrichs/Poincaré inequality and (4) is necessary to apply the unique continuation principle.
In the following, we first present a short preliminary section, which serves as a reminder for well-established concepts such as trace operators and boundary triples, which we will employ to prove our main results concerning well-posedness (Section 3) and stability (Section 4). We will end our article with a conclusion and a short appendix on the applicability of a unique continuation theorem in Section 4 and an additional result regarding regularity in the case of the trace maps taking values in $\mathrm{L}^{2}(\partial\Omega;\mathbb{C})$.
@@ 2. Preliminaries
This section is largely a short recapitulation of established theory that we present both for the convenience of the reader, as well as to establish notation that we are going to use in consecutive sections.
Note that we will often not specify the codomain of $\mathrm{L}^{p}$ spaces, Sobolev spaces and spaces of continuous functions as it will be clear from the context, e.g., we write simply $\mathrm{L}^{2}(\Omega)$ for both $\mathrm{L}^{2}(\Omega;\mathbb{C})$ and $\mathrm{L}^{2}(\Omega;\mathbb{C}^{k})$.
@@ 2.1. Sobolev Spaces
We clarify/introduce some notation first: We let $\mathrm{H}^{1}(\Omega)$ be the space of all $\mathrm{L}^{2}$-functions with distributional derivative in $\mathrm{L}^{2}(\Omega)$ also, together with the norm $\lVert\boldsymbol{\cdot}\rVert_{\mathrm{H}^{1}(\Omega)}\coloneq\sqrt{\lVert\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}^{2}+\lVert\nabla\boldsymbol{\cdot}\rVert_{\mathrm{L}^{2}(\Omega)}^{2}}$. In other words, $\mathrm{H}^{1}(\Omega)=\bigl(\operatorname{dom}(\nabla),\lVert\boldsymbol{\cdot}\rVert_{\operatorname{dom}(\nabla)}\bigr)$, where $\nabla$ is the weak gradient on $\mathrm{L}^{2}(\Omega)$. Similarly, the corresponding space for the divergence operator $\operatorname{div}f\coloneq\sum_{i=1}^{d}\partial_{i}f_{i}$ is
$\mathrm{H}(\operatorname{div},\Omega)\coloneq\big\{f\in\mathrm{L}^{2}(\Omega;\mathbb{C}^{d})\,\big|\,\mathopen{}\operatorname{div}f\in\mathrm{L}^{2}(\Omega;\mathbb{C})\ \text{(in the distributional sense)}\big\}.$
Note that $\mathrm{C}_{\mathrm{c}}^{\infty}(\overline{\Omega};\mathbb{C}^{d})\coloneq\left\{f\big|_{\overline{\Omega}}\,\middle|\,\mathopen{}f\in\mathrm{C}_{\mathrm{c}}^{\infty}(\mathbb{R}^{d};\mathbb{C}^{d})\right\}$ is dense in $\mathrm{H}(\operatorname{div},\Omega)$, see, e.g., [9, Ch. IX Part A Sec. 2 Thm. 1] or [25, Thm. 3.18].
2.2. Trace Operators
The following can be found in more detail in [18, Appendix A]. For $f\in\mathrm{C}_{\mathrm{c}}^{\infty}(\overline{\Omega})$ we can define the map
We can extend $\tilde{\gamma}_{0}$ to a continuous map $\gamma_{0}\colon\mathrm{H}^{1}(\Omega)\to\mathrm{L}^{2}(\partial\Omega)$, the Dirichlet trace operator. We define its image as
$\mathrm{H}^{\frac{1}{2}}(\partial\Omega)\coloneq\operatorname{ran}\gamma_{0},\quad\text{equipped with }\quad\lVert\phi\rVert_{\mathrm{H}^{\frac{1}{2}}(\partial\Omega)}\coloneq\inf\{\lVert g\rVert_{\mathrm{H}^{1}(\Omega)}\,|\,\mathopen{}\gamma_{0}g=\phi\}.$
This space is even a Hilbert space, since it can be represented by the quotient space $\mathrm{H}^{1}(\Omega)/\ker\gamma_{0}$. Note that $\gamma_{0}$ is a continuous map from $\mathrm{H}^{1}(\Omega)$ to $\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$ and $\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$ is continuously embedded in $\mathrm{L}^{2}(\partial\Omega)$. Let $\Gamma_{0},\tilde{\Gamma}\subseteq\partial\Omega$ be a (measurable) partition of $\partial\Omega$ (up to sets of measure zero). We can restrict $\gamma_{0}f$ to $\Gamma_{0}$, which makes $\gamma_{0}\big|_{\Gamma_{0}}\colon f\mapsto\gamma_{0}f\big|_{\Gamma_{0}}$ continuous from $\mathrm{H}^{1}(\Omega)$ to $\mathrm{L}^{2}(\Gamma_{0})$ also. Hence,
$\mathring{\mathrm{H}}^{1}_{\Gamma_{0}}(\Omega)\coloneq\ker\gamma_{0}\big|_{\Gamma_{0}}=\left\{f\in\mathrm{H}^{1}(\Omega)\,\middle|\,\mathopen{}\gamma_{0}f\big|_{\Gamma_{0}}=0\right\}$